The width or a rectangle is 2 feet less than the length. The area is 3 feet squared. Find the length and the width.
Answer by checkley77(12844)

W=L-2
A=LW
3=L(L-2)
3=L^2-2L
L^2-2L-3=0
(L]-3)(L=1)=0
L-3=0
L=3 ANS. FOR THE LENGTH.
W=3-2=1 ANS. FOR THE WIDTH.
PROOF:
3=3*1
3=3